# Budget-Adaptive Adapter Tuning in Orthogonal Subspaces for Continual Learning in LLMs

### v1

@@ 3 Experiments
We evaluate our approach using two CL benchmarks for LLMs. The first is the standard CL benchmark [31], which comprises 5 text classification datasets: AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers. The second is a continual learning benchmark consisting of a larger number of tasks with 15 datasets [20]. This benchmark includes five tasks from the standard CL benchmark, four from GLUE benchmark (MNLI, QQP, RTE, SST2) [32], five from SuperGLUE benchmark (WiC, CB, COPA, MultiRC, BoolQA) [33] and the IMDB movie reviews dataset [34]. Following the methodology of [20], we randomly select 1000 samples per class for training. The task details and training sequences of tasks used in our experiments are provided in Appendix A.5.
@@
Let $a_{i,j}$ denote the test accuracy on the $i$-th task after training on the $j$-th task. To evaluate performance, we use the mean accuracy over all tasks after completing training on the final task. Specifically, it is defined as $\frac{1}{T}\sum_{i=1}^{T}a_{i,T}$, where $T$ is the total number of tasks.
@@
We compare our method against various CL baseline approaches, including: SeqFT [35] trains all model parameters on a sequence of tasks without any regularization or replaying samples from the previous tasks. EWC [36] finetune the whole model with a regularization loss that prevents updating parameters that could interfere with previously learned tasks. LwF [37] constrains the shared representation layer to be similar to its original state before learning the new task. Inc-Adapter trains new Adapter parameters on a sequential series of tasks without any constraints or mechanism. Replay finetunes the whole model with a memory buffer, and replay samples from old tasks when learning new tasks to avoid forgetting. L2P [38] uses the input to dynamically select and update prompts from the prompt pool in an instance-wise fashion. LFPT5 [39] continuously train a soft prompt that simultaneously learns to solve the tasks and generate training samples, which are subsequently used in experience replay. O-Adapter train new Adapter parameters on a sequential series of tasks with orthogonal parameter subspace constraints. O-LoRA [27] train new LoRA parameters on a sequential series of tasks in orthogonal subspace while fixing the LoRA matrices of previous tasks. ProgPrompt [20] adopts a task-specific soft prompt for each distinct task, sequentially appending it to prior learned prompts. In essence, it trains individual models per task, leveraging the task ID to select the appropriate model during inference. PerTaskFT train a separate model for each task. MTL train a model simultaneously on all tasks as multi-task learning. This approach represents the theoretical upper bound for continual learning performance with a single model, as it maintains access to the entire task distribution throughout training, thus eliminating the fundamental forgetting challenge in continual learning scenarios. In our continual learning baseline selection, we specifically focused on methods that could be reliably reproduced to ensure fair comparison. Furthermore, to guarantee the authenticity and consistency of our experimental results, we reproduced and reran all baseline methods in our infrastructure.
@@ 3.2 Main Results
Our experiments employ the encoder-decoder T5 model [40], consistent with baselines in CL for NLP. Following previous works [39, 27], we report the results of three independent runs with different task orders on each CL benchmark, in Tab.1. All experimental results are reported as the average of three runs. For more detailed settings, refer to Appendix A.4.
Table 1: Testing performance on two standard CL benchmarks with T5-large.
Standard CL Benchmark Large Number of Tasks
Order-1 Order-2 Order-3 Average Order-4 Order-5 Order-6 Average
SeqFT 18.9 24.9 41.7 28.5 7.4 7.4 7.5 7.4
EWC 48.7 47.7 54.5 50.3 45.3 44.5 45.6 45.1
LwF 54.4 53.1 49.6 52.3 50.1 43.1 47.4 46.9
Inc-Adapter 57.5 47.8 66.1 57.1 54.3 46.1 58.1 52.8
Replay 55.2 56.9 61.3 57.8 55.0 54.6 53.1 54.2
L2P 60.3 61.7 61.1 60.7 57.5 53.8 56.9 56.1
LFPT5 68.6 72.4 76.9 72.6 69.4 67.8 68.6 68.6
O-Adapter 73.3 73.4 73.1 73.3 69.3 61.8 65.7 65.6
O-LoRA 75.0 75.4 75.6 75.3 71.2 63.8 70.6 68.7
OA-Adapter 75.7 76.2 76.1 76.0 70.9 65.2 71.4 69.2
ProgPrompt 75.1 75.0 75.2 75.1 78.0 77.7 77.9 77.9
PerTaskFT 70.0 70.0 70.0 70.0 78.1 78.1 78.1 78.1
MTL 80.0 80.0 80.0 80.0 76.5 76.5 76.5 76.5
Results on Standard Continual Learning Benchmarks.
Across all task orders of the standard CL benchmark, OA-Adapter consistently surpasses previous methods by a significant margin, as illustrated in Table 1. Notably, OA-Adapter, compared to the prior state-of-the-art method O-LoRA, achieves performance closer to MTL, the upper bound of continual learning with a single model. Additionally, the performance demonstrates a clear descending trend from OA-Adapter to O-Adapter to Inc-Adapter, providing intuitive evidence that orthogonal parameter subspace constraints effectively prevent catastrophic forgetting, while dynamic bottleneck dimension adaptation further enhances performance through efficient budget allocation. Furthermore, our approach significantly outperforms PerTaskFT, which indicates that OA-Adapter not only avoids catastrophic forgetting but also utilizes knowledge from prior tasks to enhance the learning of new tasks.
Performance with Large Number of Tasks.
On the more challenging benchmark comprising 15 tasks, OA-Adapter consistently outperforms O-LoRA, as illustrated in Table 1. While ProgPrompt shows higher accuracy, it requires task identifiers during inference and maintains separate parameters per task, fundamentally limiting its generalization to unseen tasks and practical deployment in real-world LLM applications. Notably, all continual learning methods still trail behind PerTaskFT and MTL, highlighting that continual learning for a large number of tasks remains a significant challenge.
@@
As established in Section 2.2, OA-Adapter leverages dynamic parameter budget allocation to achieve enhanced parameter efficiency. To quantify this advantage, we compare the parameter utilization between OA-Adapter and O-LoRA across various initial budget conditions, as illustrated in Table 2. For OA-Adapter, budget allocation represents bottleneck dimension distribution, while for O-LoRA, it determines the module’s intrinsic rank. Remarkably, OA-Adapter achieves superior performance while using $46.6\%$ to $58.5\%$ fewer parameters compared to O-LoRA’s fixed budget approach. Moreover, OA-Adapter maintains consistent performance excellence across all tested initial budget settings, demonstrating its robust dynamic allocation capabilities. These results highlight our method’s ability to effectively adapt parameter budget allocation according to task-specific requirements, providing substantial efficiency benefits over static parameter budget allocation approaches.
@@
Initial Budget Method Avg Final Budget Params Avg Performance
16 O-LoRA 16 4.72M 75.3
OA-Adapter 9.95 1.96M-58.5% 76.0+0.7
8 O-LoRA 8 2.36M 74.5
OA-Adapter 6.05 1.18M-50.0% 74.7+0.2
4 O-LoRA 4 1.18M 73.8
OA-Adapter 3.18 0.63M-46.6% 74.1+0.3
3.3 Discussions
Occurrence and Mitigation of Catastrophic Forgetting.
We first demonstrate the occurrence and mitigation of catastrophic forgetting over Order-1 on the standard CL benchmark. As shown in Figure 2, we observe a significant decline of each task after their training phase without orthogonal parameter subspace constraints. Performance on Task 2 deteriorates to nearly zero by the end of the subsequent two tasks, while performance on Task 1 and Task 3 significantly decreases from their respective levels at the end of their training phases. In contrast, the performance with orthogonal parameter subspace constraints remains largely preserved by the end of Task 4, with the most severe degradation limited to only $14\%$ performance loss on Task 2. These results demonstrate that severe forgetting phenomena occur during multi-task training and confirm that orthogonal parameter subspace constraints can effectively and consistently mitigate such forgetting. Similar trends are consistently observed under other task orders, as detailed in Appendix A.2.
Figure 2: Occurrence and mitigation of catastrophic forgetting during sequential training following Order-1 across multiple tasks. Solid lines represent models with orthogonal parameter subspace constraints, while dashed lines indicate models without. Each color corresponds to a specific task’s test accuracy. Colored background shading denotes the training phase for each respective task, with the X-axis scaled proportionally to accommodate varying training step durations.
Additionally, as expected, each task achieves higher performance more rapidly during its corresponding training phase without orthogonal subspace constraints, though the final performance is not substantially higher than with orthogonal parameter subspace constraints. This occurs because the model has greater flexibility in parameter update directions when not constrained to preserve previous knowledge, thus more easily finding optimal solutions. Meanwhile, this also demonstrates that while orthogonal subspace constraints limit the model’s choice of update directions, they still maintain sufficient capacity for subsequent tasks.
Interestingly, tasks exhibit brief performance recovery at the beginning of subsequent training phases before experiencing extended forgetting. Performance on Task 1 recovers to nearly $100\%$ at the start of Task 3 and to approximately $60\%$ during Task 4. Despite performance on Task 2 declining to near-zero by the end of Task 3, it shows rapid improvement during initial Task 4 training. The phenomenon resembles human memory: knowledge seemingly forgotten from disuse often requires minimal effort to reactivate. This suggests that apparent catastrophic forgetting merely masks a latent recovery potential within the model, indicating that knowledge representations remain partially preserved despite significant performance degradation.
@@ 3.3 Discussions
To assess the impact of our threshold strategy, we compare two policies: (a) a fixed, non-learnable threshold and (b) our proposed learnable threshold that adapts across different layers and tasks during training, as described in Section 2.2. These strategies are assessed using the T5-large model across three task orders on the standard continual learning benchmark. The results, depicted in Tab. 4, reveal that the dynamic threshold consistently demonstrates superior performance compared to the fixed threshold. This confirms our analysis in Section 2.2, where we demonstrate that the dynamic threshold mechanism enables bidirectional adjustment of budget allocation without introducing complex mechanisms or additional computational overhead, thereby enhancing flexibility in the optimization process.
@@
Table 4: Comparisons of model scales.
@@
We investigate the performance of models across varying model scales (T5-base, T5-large, T5-XL) using the standard continual learning benchmark. We evaluate both our method and O-LoRA across three task orders. The results, depicted in Tab. 4, reveal that OA-Adapter’s average accuracy consistently improves as the parameter size increases, which suggests that our approach effectively leverages the increased representational capacity of larger models. The consistent superiority across different scales indicates that OA-Adapter’s mechanism provides effective protection against catastrophic forgetting while enabling precise task-specific optimization, regardless of the underlying model architecture’s complexity. Moreover, OA-Adapter consistently outperforms O-LoRA across all model scales.
@@ 4 Conclusion
In this paper, we introduce OA-Adapter, a novel parameter-efficient approach for continual learning in LLMs that considers both dynamic budget adaptation and orthogonal subspace learning in a single end-to-end training stage. Our comprehensive experiments demonstrate OA-Adapter’s consistent superiority over existing methods across multiple benchmarks while using significantly fewer parameters. The observed heterogeneity in optimal parameter allocation across tasks and layers validates the necessity of our budget-adaptive approach. As the first work to integrate budget adaptation into parameter-efficient fine-tuning for continual learning in LLMs, OA-Adapter establishes a new paradigm that jointly optimizes parameter budget allocation and knowledge preservation. This advancement paves the way for more scalable, efficient, and effective adaptation of LLMs to evolving real-world applications.
@@ Appendix A Appendix
A.1 Related Work.
Continual Learning for LLMs.
Existing continual learning (CL) techniques can be broadly classified into several categories, including replay-based, regularization-based, architecture-based, optimization-based, and representation-based methods [12, 11]. Among these, replay-based, regularization-based, and architecture-based methods have been most extensively applied in the context of continual learning for LLMs. Replay-based methods [13, 14, 15] maintain a memory buffer containing data from previous tasks, which is used to retrain the model alongside new data. While these methods are valued for their simplicity, stability, and strong performance, their core operation—storing and replaying data—introduces significant privacy concerns and leads to substantial storage overheads, especially given that language datasets often contain personal or sensitive information. Regularization-based methods [16, 17, 18, 19] introduce a regularization term that penalizes significant weight deviations across tasks, thereby attempting to balance performance between old and new tasks. However, the complexity of these methods increases rapidly as the number of tasks grows, which often results in a sharp decline in performance retention on older tasks. These approaches primarily focus on learning all incremental tasks within a shared parameter space, which is a major contributor to task interference. In contrast, many architecture-based methods [20, 21, 22, 23, 24] address this challenge by proposing incorporating task-specific components, isolated parameters, or dedicated pathways within the model. These strategies essentially learn task-specific expert modules for different tasks. While these methods can effectively mitigate task interference, they often rely heavily on explicit task ID during the inference stage, which limits their ability to generalize across different tasks. Building on a similar principle, recent advancements have explored a promising direction that can retain generalization capacity while reducing task interference without requiring explicit task ID. These methods, such as OGD [25], AOP [26], and O-LoRA [27], constrain weight updates for each task to lie within mutually orthogonal subspaces of the high-dimensional parameter space. By doing so, these methods effectively decouple the optimization directions of different tasks, thereby mitigating interference and preserving performance on previously learned tasks. However, a common limitation of many existing orthogonal subspace methods is that they typically adopt a fixed budget for all tasks and layers. This uniform treatment overlooks the inherent variability in complexity and importance across different tasks and network layers, which is particularly pronounced in large LLMs. To fully exploit the potential of methods based on orthogonal subspace learning, it is essential to develop mechanisms that can dynamically adapt budget allocation based on task-specific and layer-specific requirements.
Adaptive Budget Tuning.
Although existing budget-adaptive tuning methods for LLMs differ in implementation details, they commonly adopt a multi-stage design in which budget adaptation is handled in a separate stage from model optimization. This multi-stage paradigm is often designed to identify optimal configurations for different tasks or layers in large-scale models. For example, DiffoRA [30] trains a LoRA module for each layer in the first stage, and then selects a subset of these modules based on a learned Difference-aware Adaptive Matrix in a subsequent stage. ElaLoRA [29] involves a “Warm-up” fine-tuning stage, followed by a “Dynamic Rank Adjustment” stage to adapt LoRA module ranks, and finally a “Stabilization” fine-tuning stage. AdaLoRA [28] injects rank pruning operations after training iterations by applying singular value decomposition (SVD) to selectively retain low-rank components. While these methods achieve strong performance on single-task static datasets, their multi-stage design introduces substantial computational and engineering overhead that compromises practicality and scalability, while creating misalignment between optimization objectives and budget adaptation criteria that ultimately hinders the achievement of truly optimal solutions. Moreover, their design is limited to only unidirectional rank reduction operations, which precludes bidirectional budget adaptation. These characteristics make it particularly challenging to extend such methods to the CL setting, where the training process must be efficient, unified, and adaptive to sequentially arriving tasks. Additionally, recent methods such as ALoRA [41] and ElaLoRA [29] eliminate parameters based on specific importance metrics. While these approaches offer strong interpretability, their elimination ratios rely on empirical settings rather than adapting to the intrinsic differences between sequentially arriving tasks. Imitating the formulation of AdaLoRA, SoRA [42] introduces a proximal gradient-based update rule with strong theoretical grounding. However, it relies on fixed hyperparameters to decide the fixed sparsity threshold, which renders optimization challenging. Furthermore, SoRA’s scheduling algorithm for sparsity indicators is more suitable for exploring the balance between performance and parameter efficiency after convergence in single-task scenarios, rather than capturing task sequence characteristics in multi-task environments.
A.2 Occurrence and Mitigation of Catastrophic forgetting.
To further validate the effectiveness and consistency of orthogonal parameter subspace constraints, we conduct sequential training following Order-2 and Order-3, with results illustrated in Figure 4 and 5, respectively. Consistent with the results reported in Section 3.3, we observe severe catastrophic forgetting in the absence of orthogonal constraints, especially for earlier tasks. In contrast, models trained with orthogonal parameter subspace constraints are able to preserve performance across all tasks to a much greater extent. Notably, although the specific tasks affected most by forgetting vary depending on the task order, the general trend holds: orthogonal constraints provide consistent mitigation of forgetting regardless of task permutation. These results reinforce our earlier findings and highlight the robustness of orthogonal subspace regularization as a general mechanism for alleviating forgetting in continual learning scenarios.
A.3 Heterogeneous Budget Requirements Across Tasks and Layers.
As discussed in Section 3.3, we extend our analysis to investigate budget allocation in the context of continual learning. Here, we further present results under other task orders, as illustrated in Figure 6 and 7. These findings corroborate our analysis in Section 3.3: (a) The relationship between performance and parameter budget does not follow constant rules but rather necessitates case-specific consideration. (b) Within continual learning scenarios, the first task primarily focuses on acquiring capabilities built upon the pretrained model, whereas subsequent tasks must additionally preserve knowledge from preceding tasks, thus requiring more nuanced fine-tuning. This further substantiates our analysis that different tasks in the CL scenarios require varying budgets, and that allocating budgets according to training sequence and task characteristics is both necessary and justified.
A.4 Implementation Detail.
All our experiments involving T5 models were performed on a server outfitted with four NVIDIA GeForce RTX 3090 GPUs, utilizing the DeepSpeed repository for implementation. Following previous studies [35, 43], for CL experiments, for each dataset we use the available validation set as a test set (since test data is not available) and hold out 500 samples from the train set to construct the validation set. For every sequence of tasks across different orders, we trained the models for one epoch using a batch size of 32 (8 per GPU), a dropout rate of 0.1, and no weight decay. Across all experiments, we primarily used Adapter modules with a bottleneck dimension of 16, and applied a sparsification threshold chosen from $\{1e\text{-}3,1e\text{-}4,1e\text{-}5\}$. The learning rate was selected from $\{5e\text{-}3,3e\text{-}3,1e\text{-}3,5e\text{-}4\}$ depending on task characteristics. We applied an orthogonality regularization on the Adapter’s upsampling matrix with a coefficient $\lambda_{orth}\in\{0.5,1,5\}$, and used an additional coefficient $\lambda_{2}\in\{0,0.1,0.5\}$ to scale the associated L2 loss term. This flexible configuration allowed us to balance knowledge retention and model sparsity across tasks, especially in long sequences with substantial distribution shifts. To ensure experimental comparability and fair result comparison, we maintain consistency with O-LoRA [27] by adopting instruction tuning as the training paradigm across all experiments for both our method and other baselines, as shown in Table 5. This approach offers dual advantages: it incorporates human expertise for efficient learning while enabling models to better capture underlying principles through explicit guidance, thereby enhancing generalization capabilities. The consistent instruction-based framework allows for direct performance comparisons while leveraging the benefits of natural language supervision.
Table 5: Instructions for different tasks.
Task Prompts
NLI What is the logical relationship between the "sentence 1" and the "sentence 2"?
Choose one from the option.
QQP Whether the "first sentence" and the "second sentence" have the same meaning?
Choose one from the option.
Choose one from the option.
or false? Choose one from the option.
in both sentence? Choose one from the option.
A.5 Datasets and Task Orders.

### v2

@@ 4 Experiments and Analysis
We evaluate OA-Adapter using three CL benchmarks for LLMs: 1) Standard CL Benchmark (Zhang et al., 2015): a continual learning benchmark comprising five text classification datasets: AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers. 2) Long Sequence Benchmark (Razdaibiedina et al., 2023): a continual learning benchmark of 15 classification datasets, including five tasks from the Standard CL Benchmark, four from the GLUE benchmark (MNLI, QQP, RTE, SST2) (Wang et al., 2019b), five from the SuperGLUE benchmark (WiC, CB, COPA, MultiRC, BoolQA) (Wang et al., 2019a), and the IMDB movie reviews dataset (Maas et al., 2011). 3) SuperNI Benchmark (Wang et al., 2022a): a benchmark of diverse NLP tasks with expert-written instructions, enabling rigorous benchmarking of practical continual learning settings for LLMs. Specifically, we focus on five task categories: dialogue generation, information extraction, question answering, summarization, and sentiment analysis. From each category, three tasks are selected, resulting in a sequence of 15 tasks. Following Razdaibiedina et al. (2023); Zhao et al. (2024), for all benchmarks we randomly select 1,000 samples per class for training. The task details and training sequences of tasks used in our experiments are provided in Appendix A.3.
@@
[…]
@@
We compare our method against various CL baseline approaches, including: Replay finetunes the whole model with a memory buffer, and replay samples from old tasks when learning new tasks to avoid forgetting. L2P (Wang et al., 2022b) uses the input to dynamically select and update prompts from the prompt pool in an instance-wise fashion. LFPT5 (Qin & Joty, 2021) continuously train a soft prompt that simultaneously learns to solve the tasks and generate training samples, which are subsequently used in experience replay. O-LoRA (Wang et al., 2023a) train new LoRA parameters on a sequential series of tasks in orthogonal subspace while fixing the LoRA matrices of previous tasks. ProgPrompt (Razdaibiedina et al., 2023) adopts a task-specific soft prompt for each distinct task, sequentially appending it to prior learned prompts. In essence, it trains individual models per task, leveraging the task ID to select the appropriate model during inference. In our continual learning baseline selection, we specifically focused on methods that could be reliably reproduced to ensure fair comparison. Furthermore, to guarantee the authenticity and consistency of our experimental results, we reproduced all baseline methods in our infrastructure.
@@ 4.2 Main Results
Our experiments employ both the encoder-decoder T5 model (Raffel et al., 2020) and the decoder-only LLaMA-7B model (Touvron et al., 2023), consistent with baselines in CL for NLP. Following previous works (Qin & Joty, 2021; Wang et al., 2023a), we report the results of three independent runs with different task orders on each CL benchmark, in Table 1. All experimental results are reported as the average of three runs. For more detailed settings, refer to Appendix A.4.
Table 1: Testing performance on three benchmarks with T5-large.
Standard CL Benchmark Long Sequence Benchmark SuperNI Benchmark
AA FWT BWT AA FWT BWT AA FWT BWT
Replay 57.8 -8.4 -13.4 54.2 1.2 -12.2 20.5 -1.4 -15.8
L2P 60.7 -3.1 -11.2 56.1 1.4 -16.6 12.7 -19.1 -8.0
LFPT5 72.6 -1.9 -8.3 68.6 2.7 -12.8 26.7 -0.5 -14.5
O-LoRA 75.3 -3.6 -9.1 68.7 -6.2 -4.1 25.9 -7.8 -24.6
ProgPrompt 75.1 -2.3 -8.1 63.2 -2.3 -6.7 23.3 -3.3 -13.2
OA-Adapter 76.0 -2.7 -7.5 69.2 -4.4 -3.2 29.3 -5.2 -6.0
Results on Continual Learning Benchmarks.
As illustrated in Table 1, across all task orders of the standard CL benchmark, OA-Adapter consistently surpasses previous methods by a significant margin. On the other two benchmarks comprising 15 tasks, OA-Adapter consistently outperforms previous methods. Notably, on the SuperNI dataset, which is primarily designed for generative tasks, CL methods still perform poorly, underscoring that continual learning for complex tasks remains a significant challenge. To disentangle the roles of orthogonal constraints and adaptive budget, we compare the full OA-Adapter with two variants: a no-orthogonality variant that removes the orthogonal parameter subspace constraint, and a fixed budget variant that disables dynamic bottleneck dimension adaptation. As shown in Table 3, removing orthogonality leads to significant degradation, underscoring its importance in mitigating catastrophic forgetting; disabling adaptation also reduces performance, indicating that adaptive capacity allocation provides additional gains.
@@
As established in Section 3.2, OA-Adapter leverages dynamic parameter budget allocation to achieve enhanced parameter efficiency. We compare the parameter utilization between OA-Adapter and O-LoRA across various initial budget conditions, as illustrated in Table 3. For OA-Adapter, budget allocation represents bottleneck dimension distribution, while for O-LoRA, it determines the module’s intrinsic rank. Remarkably, OA-Adapter achieves superior performance while using $46.6\%$ to $58.5\%$ fewer parameters compared to O-LoRA’s fixed budget approach. Moreover, OA-Adapter maintains consistent performance excellence across all tested initial budget settings, demonstrating its robust dynamic allocation capabilities. These results highlight our method’s ability to effectively adapt parameter budget allocation according to task-specific requirements, providing substantial efficiency benefits over static parameter budget allocation approaches.
Table 2: Ablation study evaluating the effects of orthogonality and budget adaptation.
Method Dataset
Standard Long SuperNI
OA-Adapter 76.0 69.2 29.3
- Orthogonality 57.1 52.8 13.7
- Budget Adaptation 73.3 65.6 24.2
@@
Method Budget Allocation Params AA
O-LoRA 16→16 4.72M 75.3
OA-Adapter 16→9.95 1.96M-58.5% 76.0+0.7
O-LoRA 8→8 2.36M 74.5
OA-Adapter 8→6.05 1.18M-50.0% 74.7+0.2
O-LoRA 4→4 1.18M 73.8
OA-Adapter 4→3.18 0.63M-46.6% 74.1+0.3
@@
4.3 Discussions and Analysis
Figure 3: Occurrence and mitigation of catastrophic forgetting during sequential training following Order-1 across multiple tasks.
Occurrence and Mitigation of Catastrophic Forgetting.
As shown in Figure 3, we demonstrate the occurrence and mitigation of catastrophic forgetting on the standard CL benchmark. Solid and dashed lines denote models with and without orthogonal parameter-subspace constraints, and each color corresponds to a distinct task. Shaded background bands mark the training phase of each task, and the X-axis is scaled proportionally to training steps. We observe a significant decline of each task after their training phase without orthogonal constraints. Performance on Task 2 drops to nearly zero by the end of the subsequent tasks, while Task 1 and Task 3 also decrease substantially. In contrast, the performance with orthogonal constraints remains largely maintained through Task 4, with the most severe degradation limited to only $14\%$ performance loss on Task 2. These results demonstrate that severe forgetting phenomena occur during multi-task training and orthogonal parameter subspace constraints can effectively mitigate it. Similar trends are consistently observed under other task orders, as detailed in Appendix A.6. Additionally, as expected, each task achieves higher performance more rapidly during its corresponding training phase without orthogonal constraints, though final performance is not substantially higher than with orthogonal constraints. This suggests that the model has greater flexibility in parameter update directions when not constrained to preserve previous knowledge, and orthogonal subspace still maintain sufficient capacity for subsequent tasks. Interestingly, tasks exhibit brief performance recovery at the beginning of subsequent training phases before further decline. Performance on Task 1 recovers to nearly $100\%$ at the start of Task 3 and to approximately $60\%$ during Task 4. Although Task 2 drops to near-zero by the end of Task 3, it improves rapidly early in Task 4 training. This behavior resembles human memory, where disused knowledge can be reactivated with minimal effort. This suggests that catastrophic forgetting may only superficially mask a latent capacity, and knowledge representations remain partially preserved despite significant performance degradation.
@@ 4.3 Discussions and Analysis
To assess the impact of our threshold strategy, we compare two policies: (a) a fixed, non-learnable threshold and (b) our proposed learnable threshold that adapts across different layers and tasks during training, as described in Section 3.2. These strategies are assessed using the T5-large model across three task orders on the standard continual learning benchmark. The results, depicted in Table 5, reveal that the dynamic threshold consistently demonstrates superior performance compared to the fixed threshold. This confirms that the dynamic threshold mechanism enables bidirectional adjustment of budget allocation without introducing complex mechanisms, thereby enhancing more flexibility in the optimization process.
@@
Table 5: Comparisons of pre-trained models.
@@
We investigate the performance of OA-Adapter and O-LoRA across varying models (T5-base, T5-large, T5-XL, Llama-7B) using the standard continual learning benchmark. The results, depicted in Table 5, reveal that OA-Adapter’s average accuracy consistently improves, which suggests that OA-Adapter effectively leverages the increased representational capacity of larger models. The consistent superiority across different models indicates that OA-Adapter’s mechanism provides effective protection against catastrophic forgetting while enabling precise task-specific optimization, regardless of the underlying model architecture’s complexity and scales.
Compatibility with Replay.
To assess compatibility with replay-based strategies, we evaluate a hybrid variant OA-Adapter+Replay and compare it with the strong replay baseline SAPT (Zhao et al., 2024). The hybrid consistently yields additional gains while preserving low forgetting, indicating that our method complements replay. Full results and settings are provided in Appendix A.7.
@@ 5 Conclusion
In this paper, we introduce OA-Adapter, a novel parameter-efficient approach for continual learning in LLMs that considers both dynamic budget adaptation and orthogonal subspace learning in an end-to-end training stage. Our comprehensive experiments demonstrate OA-Adapter’s consistent superiority over existing methods across multiple benchmarks while using significantly fewer parameters. The observed heterogeneity in optimal parameter allocation across tasks and layers validates the necessity of our budget-adaptive approach. As the first work to integrate budget adaptation into parameter-efficient fine-tuning for continual learning in LLMs, OA-Adapter establishes a new paradigm that jointly optimizes parameter budget allocation and knowledge preservation. This advancement paves the way for efficient and effective adaptation of LLMs to evolving real-world.
6 Reproducibility Statement
It is important to note that the work presented in this paper is reproducible. To ensure the reproducibility of our results, we have made several efforts, which we summarize below. First, the experiments conducted in this paper, including the models and datasets used, are described in detail in Section 4. For more specific information on the datasets, we refer readers to Appendix A.3, which provides the complete details. The implementation details of the models, including hyperparameter settings and optimization procedures, are provided in Appendix A.4. We have submitted the source code as supplementary material, which includes all implementations necessary to reproduce the experiments. After acceptance, the code will be publicly available on GitHub for the research community. This code includes the OA-Adapter model, training routines, and configuration files. By providing these detailed resources, we aim to ensure that our work can be reproduced accurately. Furthermore, we encourage others to conduct further exploration and research based on our work.
@@ Appendix A Appendix
A.1 LLM Usage Declaration
Large language models (e.g., ChatGPT) were used solely for language editing and formatting. They did not contribute to the conception, design, implementation, analysis, data generation or labeling, or evaluation of the methods and results. All technical content and claims were authored and verified by the authors, and no personal, proprietary, or sensitive data were shared with LLM services.
A.2 Details About Orthogonality Constraints
In Equation (8) and (9), OA-Adapter choose $\mathcal{W}_{2}^{(k)}\cdot\Gamma^{(k)}$ rather than $\Gamma^{(k)}\cdot\mathcal{W}_{1}^{(k)}$ to enforce the orthogonal relationships. Here’s why:
[…]
Furthermore, in the full OA-Adapter, we introduce a diagonal gate $\Gamma^{(k)}$, for each task $k$ to selectively activate a subset of the basis directions, and then we impose the orthogonality constraint on the product $W_{2}^{(k)}\Gamma^{(k)}$. This choice is deliberate: $W_{2}^{(k)}\Gamma^{(k)}$ directly determines the directions injected into the transformer’s hidden state, whereas $W_{1}^{(k)}$ only controls how strongly each direction is used. If orthogonality were enforced on $\Gamma^{(k)}W_{1}^{(k)}$ instead, the subsequent multiplication by $W_{2}^{(k)}$ would generally break that orthogonality, allowing task interference to resurface. Constraining $W_{2}^{(k)}\Gamma^{(k)}$ keeps the basis directions themselves orthogonal across tasks, which is precisely what reduces interference in continual learning.
A.3 Datasets and Task Orders.